Express the following in rational form
a) 0.3434343434343434343434...
b)2.457777777777777777777777...

Answer by Alan3354(69443)

Express the following in rational form
a) 0.3434343434343434343434...
---
x = 0.3434343434343434343434...
100x=34.34343434343434343434...
- x = 0.3434343434343434343434...
---------------------------------------
99x= 34
x = 34/99
----------------
b)2.457777777777777777777777...
---
x = 2.457777777777777777777777... = 2.45 + 0.00777777777777...
y = 0.00777777777777...
10y =0.0777777777777...
-y = 0.00777777777777...
---------------------------
9y = 0.07
y = 0.07/9 = 7/900
2.457777777777777777777777... = 2.45 + 7/900 = 2 + 405/900 + 7/900 = 2212/900
= 553/225